OpenStudy (anonymous):

PLEAZE HELP !!!!!!! A system of equations is shown below: 5x - 5y = 10 3x - 2y = 2 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. Part B: Show that the equivalent system has the same solution as the original system of equations.

OpenStudy (anonymous):

@ganeshie8 @Hero @chmvijay @paki @tkhunny @zepdrix @nincompoop @Vocaloid @rvc @Kitten_is_back @dan815

OpenStudy (anonymous):

I will try to solve your system of equations. 5x−5y=10;3x−2y=2 Step: Solve5x−5y=10for x: 5x−5y+5y=10+5y(Add 5y to both sides) 5x=5y+10 5x 5 = 5y+10 5 (Divide both sides by 5) x=y+2 Step: Substitutey+2forxin3x−2y=2: 3x−2y=2 3(y+2)−2y=2 y+6=2(Simplify both sides of the equation) y+6+−6=2+−6(Add -6 to both sides) y=−4 Step: Substitute−4foryinx=y+2: x=y+2 x=−4+2 x=−2(Simplify both sides of the equation) Answer: x=−2 and y=−4 I hope this helps. (:

OpenStudy (anonymous):

Great work @x3mbD4rkSt4rx , you are super-duper fast.

OpenStudy (anonymous):

Thanks lol xD. I had this problem awhile ago with helping my sister. So, I had it in my notes on my computer. So anytime!(:

OpenStudy (anonymous):

And that's what OpenStudy is. "Learning goes viral, on openstudy".

OpenStudy (anonymous):

Indeed it does.

OpenStudy (mrnood):

@x3 @2know However - you r solution is NOT what the question demands You have solved by substitution but the question explicitly asks for an elimination method @layla.McCoy Are you still looking at this?

OpenStudy (anonymous):

@MrNood

OpenStudy (tkhunny):

1) Don't post problems with none of YOUR work showing. 2) Certainly don't tag the world unless you FIRST show YOUR work.

OpenStudy (anonymous):

eh i just answered. For assistance.

OpenStudy (bubblegum.):

Its better than doing nothing in this situation rn